II. Compute the escape speed for an electron from the surface of a uniformly charged sphere of radius 1.22cm and total charge 1.76x10-15 C. Neglect gravitational forces.


1
Expert's answer
2021-05-18T11:09:37-0400

The potential at the surface of the sphere is


"V=\frac{Q}{4\pi\epsilon_0 R}."

The work required to remove the electron is


"W=Vq=\frac{Qe}{4\pi\epsilon_0R}."


To escape from the sphere, the kinetic energy must be greater than this value (work-energy theorem):


"E_k=\frac12mv^2=E_p,\\\space\\
\frac12mv^2=\frac{Qe}{4\pi\epsilon_0R},\\\space\\
v=\sqrt{\frac{Qe}{2\pi\epsilon_0mR}}=21.4\text{ km/s}."